OpenStudy (anonymous):

Given a triangle with the measure of the first angle is 7x, the measure of the second angle is (x + 6) and the measure of the third angle is (2x + 4), find the measures in degrees of all three angles.

OpenStudy (anonymous):

because the interior degrees of a triangle add up to 180 degrees, you can say 7x+(x+6)+(2x+4)=180 because addition is commutative, meaning you can add in any order, you can rewrite it combining like terms 10x+10=180 simplify. 10x=170 x=17 Now, you can plug x into the expression for each angle and find the value of the angle.

OpenStudy (anonymous):

7(17)=119 (17)+6=23 2(17)+4=38 119+23+38=180